OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#4287 — LCL Bruxelles
Incident Report for Network & Infrastructure
Resolved
We are facing a problem on one of our equipments at Brussels Interxion. This impacts the traffic towards the BNIX and ChtIX customers connected to this site.


Update(s):

Date: 2010-06-16 11:11:58 UTC
We are back to normal. The chtix customers on LCL are now restored.
In case of problem, please do not hesitate to contact noc@ovh.net.

Date: 2010-06-16 10:54:05 UTC
We reswitch the infra on the LCL in the nominal config.

Date: 2010-06-16 10:49:48 UTC
So, it seems that the problem is due to a loop on BNIX.
Our switches were overloaded and did not respond to pings. We
therefore thought about a hardware from our side.

We have re-put in place all of the BNIX peers. There is
no other problem..

Date: 2010-06-16 10:45:18 UTC
Jun 16 11:15:41 20G.bru-1-6k.routers.ovh.net 147: Jun 16 10:15:23 GMT: %IPV6-4-DUPLICATE: Duplicate address FE80::213:5FFF:FE1C:6500 on TenGigabitEthernet1/2.20
Jun 16 11:05:11 20G.bru-1-6k.routers.ovh.net 88: .Jun 16 10:04:55 GMT: %IPV6-4-DUPLICATE: Duplicate address FE80::213:5FFF:FE1C:6500 on TenGigabitEthernet1/2.20
Jun 16 09:33:42 20G.bru-1-6k.routers.ovh.net 606896: Jun 16 08:33:25 GMT: %IPV6-4-DUPLICATE: Duplicate address FE80::213:5FFF:FE1C:6500 on TenGigabitEthernet1/2.20
Jun 16 09:26:29 20G.bru-1-6k.routers.ovh.net 606853: Jun 16 08:26:12 GMT: %IPV6-4-DUPLICATE: Duplicate address FE80::213:5FFF:FE1C:6500 on TenGigabitEthernet1/2.20


Date: 2010-06-16 10:45:08 UTC
Hi,

There's currently an issue on the BNIX. We're investigating and hope to
resolve it as soon as possible.

Best regards,
Dear Member,

Aside from the first looped port, the BNIX staff identified a second
looped port and removed this one ASAP from the config. Everything is
back to normal now.

We acknowledge that the error lies with BNIX and we wish to apologize
for the degraded service.

Kind Regards,

-- BNIX Tech Team

Date: 2010-06-16 10:44:42 UTC
We have a loss of packets on the 6K / F10 link.
We have updated the bru-1-6k router
http://travaux.ovh.com/?do=details&id=4288
always the same.

We thank that the optics which is in the router
is not really good for the link between both data centres.
It functioned with a switch between both of them, yet, it
did not work directly.

We will be on the scene to replace the switch
which was broken today by a spare.

Around 12h30, the problem must be resolved.

Meanwhile, we will cut off BNIX.

Date: 2010-06-16 08:59:53 UTC
To circle the problem temporarily, we bypassed bru-2-f10 in order to restore the traffic to bru-3 (Interxion) and BNIX. Only chtix customers
on LCL will be still affected.


Date: 2010-06-16 08:54:48 UTC
The problem comes probably from the link bru-1<>bru-2 on LCL. This can be a default as well on the switch force10 itself. We will send a technician
to the scene with all the necessary equipment.

Date: 2010-06-16 08:19:59 UTC
It is a problem on a switch situated in upstream at the level of LCL Diegem. We intend to make a restarting of the equipment. If it fails,
we will send a technician in emergency for a replacement with a spare.
Posted Jun 16, 2010 - 07:03 UTC